the sum of two numbers is 16. one number is 4 less than 3 times the other. find the numbers
Answer by harpazo(655)

x + y = 16
x = 3y - 4
===========


3y - 4 + y = 16
4y - 4 = 16
4y = 16 + 4
4y = 20
y = 20/4
y = 5.
We now know that y = 5.
Let us plug y = 5 into either original equation to find x.
I will use x + y = 16.
x + 5 = 16
x = 16 - 5
x = 11
The two numbers are 5 and 11.